An inverted regular square pyramid has an altitude of 12m and with top edge of 3m is fully–filled with water What is the volume of the water in m3? A portion of water is to be removed so that water will be 7m deep.
1
Expert's answer
2020-12-03T19:59:49-0500

V = (1/3)*A*h

A = area of Base = b² = 3² = 9m²

h = altitude = 12m

V = (1/3)*9*12

V = 36m³

The volume of water when it's fully filled in the cylinder is 36m³.


If a portion of water is to be removed so that water will be 7m deep:

Final height, h = 7m

Initial height, H = 12m

Final base edge, b = bm

Initial base edge, B = 3m

(H/h) = (B/b)

(12/7) = (3/b)

Solving b = 1.75m

A = b² = 1.75² = 3.0625m²

V =(1/3)*A*h

V =(1/3)*3.0625*7

V = 7.146m³

The volume of the water in the cylinder when a portion is removed so that the water is 7m deep is 7.146m³.